OpenStudy (anonymous):

PLEASE HELP the approximate value of y=sqrt(3+e^x) at x=0.08, obtained from the tangent to the graph at x=0, is 1) 2.01 2) 2.02 3) 2.03 4) 2.04 5) 2.05 explanation would be greatly appreciated

OpenStudy (mathmale):

this is "linear approximation." Have you used this method before? You are asked to approximate the value of y=sqrt(3+e^x) at x=0.08. Let y = f(x). Then the relevant linear approx. formula is \[f(x)approxf(x _{0})+f'(x _{0})*(x-x _{0})\]

OpenStudy (mathmale):

Here, f(x) = Sqrt(3+e^x) and x_0 is 0.08.

OpenStudy (anonymous):

i have never used this method by i've heard of "linear approximations"

OpenStudy (mathmale):

What method is your teacher (or your online instruction) using for linear approximations? Note that you are asked to approximate the value of the given function at x=0.08 using the tangent line. The formula I have given you does exactly that. \[f'(x _{0})\

OpenStudy (mathmale):



OpenStudy (anonymous):

I would approximate x=0.08 to x\[\approx\]0

OpenStudy (anonymous):

and solve with a point slope formula

OpenStudy (mathmale):

No, Caden. You are asked to approximate the value of the given function when x=0.08. If you want to use the point-slope form of the equation of a straight line, you'll still need to determine the slope of the line; how would you do that?

OpenStudy (anonymous):

i'm not sure that was just the first thought i had to answering this question

OpenStudy (mathmale):

What is the name of the TOPIC you are studying right now? I call it "linear approximations." But then I realize you probably haven't studied calculus yet. What is the name of your math course?

OpenStudy (anonymous):

ap calculus ab

OpenStudy (anonymous):

I'm doing a practice problem and i got extremely confused so i came here

OpenStudy (mathmale):

advanced placement calculus. Wow. congrats. But this means you are responsible for topics such as "linear approximations." \ If you give me a function and a particular x value (such as 0.08)

OpenStudy (anonymous):

are linear approximations something i should have learned before calculus?

OpenStudy (mathmale):

No, Caden, but rather "linear approximations" are an application of calculus in your first course in the subject. Perhaps your teacher is using a different name for this.